OpenStudy (anonymous):

2tan37.tan53 - sec65/cosec25 + tan^2(45) @Pawanyadav

OpenStudy (anonymous):

Show the steps.

OpenStudy (anonymous):

??? I'm confused :(

OpenStudy (pawanyadav):

{2tan37tan65=2} {tan^2(45)=1} {sec65/cosec25=sin25/cos65=sin(90-65)/cos65 =cos65/cos65=1} Now put all three values in question.

OpenStudy (anonymous):

It's 2tan37.tan53 and not 2tan37.tan65

OpenStudy (pawanyadav):

Yes, my mistake.sorry Do you understand now

OpenStudy (anonymous):

No :(

OpenStudy (pawanyadav):

Which part you not understand.

OpenStudy (anonymous):

The whole thing.

OpenStudy (anonymous):

2tan37.tan53=2 sec65/cosec25=1 tan^2(45)=(TAN45)^2=1 S0, 2-1+1 =2

OpenStudy (anonymous):

How is 2tan37.tan53 = 2?

OpenStudy (anonymous):

YOU ASKED THAT QUESTION IN UR RECENT POST

OpenStudy (anonymous):

Yeah.Okay,but then my net went off for DAYS so............

OpenStudy (pawanyadav):

2×3/4×4/3=2 In this 3/4 And4/3 are values of tan37 and tan53

OpenStudy (anonymous):

And Thanks Yaar @Pawanyadav

OpenStudy (anonymous):

SO HOPE U NOW UNDERSTAND ?

OpenStudy (anonymous):

Why are you yelling?

OpenStudy (anonymous):

HMMMM.SORRY . IT WAS JUST A QUESTION TO SEE IF YOU UNDERSTAND NOW